### v1

# Complexity of Einstein-Maxwell-non-minimal coupling $R^{2}F^{2}$: the role of the penalty factor

Abstract
In this paper, we examine the holographic complexity of a theory of Einstein-Maxwell with a non-minimal coupling $R^{2}F_{\mu\nu}F^{\mu\nu}$ term via complexity=anything. We introduce a perturbative black brane solution in AdS spacetime up to the first order of the non-minimal coupling. Because the model exhibits the resistivity proportional to the temperature, it could describe strange metals. The complexity of the solution could be characterized by three quantities: the conserved charge, the non-minimal coupling and the generalized term in the complexity. The generalized term is chosen to be $C^{2}$, the Weyl squared tensor, $R^{2}F^{2}$ and $F^{2}$. The physical interpretation of the three quantities has been provided where the generalized term is interpreted as a bulk penalty factor, the non-minimal coupling and the conserved charge as an effective scrambling time of the dual theory which contains the choice of the generalization parameter.
Keywords: AdS/CFT duality, Complexity, Black brane, non-minimal coupling, superconducting circuit, penalty factor
1 Introduction
The AdS/CFT (Anti-de Sitter/Conformal Field Theory) duality provides a powerful framework for understanding the relationship between gravitational theories in asymptotically AdS spaces and their corresponding quantum field theories on the boundary [1]. Although, our universe is a dS space, studying the duality in AdS is valuable to find a deep insights to develop the duality to dS/CFT and real-world physics [2, 3, 4, 5, 6, 7, 8]. This duality suggests that complex gravitational phenomena can be analyzed through the lens of quantum information theory, enabling the exploration of measures such as the quantum computational complexity. The complexity can be viewed as a measure of how difficult it is to prepare a given quantum state from a reference state using a sequence of operations [9]. The complexity serves as a crucial indicator of the degrees of freedom [10, 11] and the scrambling dynamics in a quantum system, revealing insights into how information is processed and stored in black hole backgrounds. By examining complexity within the AdS/CFT framework, we gain valuable insights into the structure of quantum states and their holographic properties, ultimately bridging the gap between deep gravitational insights and operational aspects of quantum field theories.
Recent developments in theoretical physics have established intriguing relationships between quantum complexity and fundamental physical quantities, notably through the lens of the Einstein-Rosen bridge volume. It suggests that the growth of the Einstein-Rosen Bridge (ERB) in an eternal AdS black hole can be explained by ER=EPR. It is given by the maximum volume of a hypersurface (ERB) anchored in the boundary time slice $\Sigma_{CFT}$, where the boundary CFT state live [12]:
$\mathcal{C}(\Sigma_{CFT})=\max_{\partial\Sigma=\Sigma_{CFT}}\frac{V(\partial\Sigma)}{G_{N}L},$ (1)
here $\Sigma$ is extremal hypersurface anchored in the boundary time slice $\Sigma_{CFT}$. On the other hand, the other conjecture that relates complexity to the action of the gravitational theory restricted to the Wheeler-DeWitt (WDW) patch, expressed as [13]:
$\mathcal{C}(\Sigma)=\frac{I_{WDW}}{\pi\hbar}.$ (2)
In [14], it is defined the spacetime volume for the WDW patch as:
$\mathcal{C}(\Sigma)=\frac{V_{WDW}}{G_{N}L^{2}}.$ (3)
Recently, it has been proposed that there is a generalization in the complexity and the volume of an arbitrary region of spacetime [15].
The exploration of non-minimal theories, those that connect the gravitational field with other fields through cross terms involving the curvature tensor, has gained traction as alternative gravity theories over time, particularly the non-minimal coupling has a significance on high curvature spacetimes. Non-minimal field theories can be categorized into five classes, based on the types of fields that non-minimally interact with gravitation:
1) The first category involves the coupling of scalar fields to spacetime curvature, originating from the Scherrer-Jordan-Thiry-Brans-Dicke theory, where a scalar field interacts non-minimally with the Ricci scalar. Various authors [16, 17, 18, 19] had distinct motivations for proposing this theory; for a historical review, see [20]. Additionally, a scalar field conformally coupled to gravity was introduced in [21], with a review provided in [22].
2) The second category focuses on modeling non-minimal interactions between the electromagnetic field and curvature, often referred to as the non-minimal Einstein-Maxwell model (see, for example, [23, 24, 25, 26]).
3) The third class includes Einstein-Yang-Mills models characterized by SU(N) symmetry [27, 28].
4) The fourth category consists of Einstein-Yang-Mills-Higgs models [29, 30].
5) Lastly, the fifth class encompasses models featuring an axion pseudo-scalar field, which couples to both electromagnetic and gravitational fields, known as non-minimal Einstein-Maxwell-axion models [31].
Non-minimal coupling in gravity, particularly when invoking the AdS (Anti-de Sitter) framework, has gained significant attention due to its potential to enhance our understanding of gravitational dynamics in various physical contexts. Due to the curvature coupling, these non-minimal theories have a significant effect on the higher curvature regimes such as near black hole or the early universe. The non-minimal term could be viewed as deforming the boundary theory, which means that n-point functions and transport coefficients would be altered. Specifically, this term breaks the translational symmetry which is more appropriate to model realistic crystals where containing impurity [32]. In addition, the non-minimal theories could model strange metals where the resistivity scales linearly with temperature [33]. The model we have considered in this paper represents a linear relation between the resistivity and the temperature which signaling the behavior of strange metals [34].
The organization of the paper is as follows: we introduce a perturbative solution for a non-minimal coupling $R^{2}F^{2}$ of Einstein-Maxwell theory in the section 2. Subsequently, the complexity growth rate of the solution in the complexity equals anything proposal is considered in the section 3. The complexity growth rate in this solution is dependent to the conserved charge $Q$, the non-minimal coupling $q_{2}$ and the term of the generalization $\gamma b$. There are some choices for the generalization of the complexity where we have chosen the Weyl squared tensor $\gamma C^{2}$, $\gamma F^{2}$ and $\gamma R^{2}F^{2}$. Then, in the section 4 we provide the physical interpretations of the conserved charge as a kind of restriction on the simple unitary operators, the non-minimal coupling as the effective scrambling time of the quantum circuit and the choice of the generalization in the quantum circuit as the penalty factor in the quantum circuit. At the end of the section 4 we provide two examples in superconducting circuits that qualitatively fit with the penalty factor interpretation and finally we conclude in the section 5.
2 Non-minimal $R^{2}F_{\mu\nu}F^{\mu\nu}$ AdS Black Brane Solution
The non-minimal Einstein-Maxwell theory, featuring a negative cosmological constant, is represented by the following action [35],[36],[37]:
$S=\int d^{4}x\sqrt{-g}\bigg[\frac{1}{\kappa}(R-2\Lambda)-\frac{q_{1}}{4}F_{\mu\nu}F^{\mu\nu}+q_{2}R^{2}F_{\mu\nu}F^{\mu\nu}\bigg],$ (4)
where $\kappa$ is the gravitational constant, $R$ is the Ricci scalar, $\Lambda=-\frac{3}{L^{2}}$ is the cosmological constant, $L$ representing the AdS radius. The parameter $q_{1}$ is dimensionless.
The Maxwell field strength $F_{\mu\nu}$ is given by:
$F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu},$ (5)
with the gauge coupling constant set to 1 and $A_{\nu}$ representing the Maxwell potentials. The parameter $q_{2}$ is a dimensionful coupling constant which also acts as the interaction term between the gauge field and the Ricci scalar.
Varying the action (4) with respect to the spacetime metric $g_{\mu\nu}$ yields the field equations:
$R_{\mu\nu}-\tfrac{1}{2}g_{\mu\nu}R+\Lambda g_{\mu\nu}=\kappa T^{\text{(eff)}}_{\mu\nu}$ (6)
where,
$T^{\text{(eff)}}_{\mu\nu}=q_{1}T^{\text{(M)}}_{\mu\nu}+q_{2}T^{(I)}_{\mu\nu}$ (7)
The energy-momentum tensor for the Maxwell field is given by,
$T^{\text{(M)}}_{\mu\nu}=\tfrac{1}{4}g_{\mu\nu}F_{\alpha\beta}F^{\alpha\beta}-F_{\mu}{}^{\alpha}F_{\nu\alpha},$ (8)
and the additional interaction term is,
$T^{(I)}_{\mu\nu}=2F_{\alpha\beta}F^{\alpha\beta}R_{\mu\nu}R+2F_{\mu}^{\alpha}F_{\nu\alpha}R^{2}-\tfrac{1}{2}F_{\alpha\beta}F^{\alpha\beta}g_{\mu\nu}R^{2}$
$+8F^{\beta\gamma}g_{\mu\nu}\nabla_{\alpha}F_{\beta\gamma}\nabla^{\alpha}R+4F^{\alpha\beta}g_{\mu\nu}R\nabla_{\gamma}\nabla^{\gamma}F_{\alpha\beta}$
$+2F_{\alpha\beta}F^{\alpha\beta}g_{\mu\nu}\nabla_{\gamma}\nabla^{\gamma}R+4g_{\mu\nu}R\nabla_{\gamma}F^{\alpha\beta}\nabla^{\gamma}F_{\alpha\beta}$
$-2F^{\alpha\beta}R\nabla_{\mu}\nabla_{\nu}F_{\alpha\beta}-4R\nabla_{\mu}F^{\alpha\beta}\nabla_{\nu}F_{\alpha\beta}-4F^{\alpha\beta}\nabla_{\mu}R\nabla_{\nu}F_{\alpha\beta}$
$-4F^{\alpha\beta}\nabla_{\mu}F_{\alpha\beta}\nabla_{\nu\ }R-2F^{\alpha\beta}R\nabla_{\nu}\nabla_{\mu}F_{\alpha\beta}-2F_{\alpha\beta}F^{\alpha\beta}\nabla_{\nu}\nabla_{\mu}R.$ (9)
Varying the action (4) with respect to $A_{\mu}$ results in the field equations:
$\nabla_{\mu}\Big(-\frac{q_{1}}{2}F^{\mu\nu}+2q_{2}F^{\mu\nu}R^{2}\Big)=0.$ (10)
Considering the equations of motion for the field $A^{\mu}$ reveals that the sign of the coupling constant $q_{2}$ should be negative. The coefficient of the kinetic term would be:
$-\frac{1}{4}\Big(q_{1}-4q_{2}R^{2}\Big),$ (11)
at large curvature regime, when $4q_{2}R^{2}\gg q_{1}$ the coefficient of the kinetic term in the Maxwell equation flips sign which means the ghost instability. Then, the sign of the coupling constant $q_{2}$ should be negative. Considering that the spacetime in our model is 4-dimensional with a two-dimensional spatial direction, we adopt the following ansatz for the metric:
$ds^{2}=-e^{-2H(r)}f(r)dt^{2}+\frac{dr^{2}}{f(r)}+\frac{r^{2}}{L^{2}}(dx^{2}+dy^{2}).$ (12)
We will use this ansatz to solve Eq. (10), where the potential 1-form is expressed as:
$A_{\mu}=\left(h(r),0,0,0\right).$ (13)
Now, we can formulate the equations of motion based on Eqs. (10) - (13) (see Appendix A).
The solution of $h(r)$ is as:
$h(r)=C_{1}\int^{r}\frac{e^{-H(u)}u^{2}}{B(u)}du+C_{2},$ (14)
$B(u)=4q_{2}\left(u\left(f^{\prime}\left(4-3uH^{\prime}\right)+uf^{\prime\prime}\right)+2f\left(\left(uH^{\prime}-1\right)^{2}-u^{2}H^{\prime\prime}\right)\right)^{2}-q_{1}u^{4}$ (15)
We will solve the metric and the gauge field equations perturbatively up to first order in $q_{2}$ and consider the following forms for $f(r)$, $H(r)$ and $h(r)$:
$f(r)=f_{0}(r)+q_{2}f_{1}(r),$ (16)
$h(r)=h_{0}(r)+q_{2}h_{1}(r),$ (17)
$H(r)=H_{0}(r)+q_{2}H_{1}(r),$ (18)
where $f_{0}(r)$, $h_{0}(r)$ and $H_{0}(r)$ are the leading order solutions of Einstein-Maxwell AdS black brane in four dimensions.
The $h_{0}(r)$ , $f_{0}(r)$ and $H_{0}(r)$ are found exactly as,
$h_{0}(r)=C_{2}-C_{1}\int^{r}\frac{1}{q_{1}u^{2}}du=Q(\frac{1}{r}-\frac{1}{r_{h}}),$ (19)
where $C_{1}=q_{1}Q$ and $C_{1}=-C_{2}$,
$f_{0}(r)=-\frac{2m_{0}}{r}-\frac{\Lambda r^{2}}{3}-\frac{\kappa q_{1}}{r}\int^{r}u^{2}h_{0}^{\prime 2}du=-\frac{2m_{0}}{r}+\frac{r^{2}}{L^{2}}+\frac{\kappa q_{1}Q^{2}}{r^{2}},$ (20)
Since the equations for the tttt–component and the $rr$-component are identical up to first order, we conclude that $8f_{0}(r)r^{3}H_{0}^{\prime}(r)=0$. Therefore, we determine that $H_{0}$ is given as follows:
$H_{0}(r)=0.$ (21)
The blackening factor at the event horizon must be suppressed, i.e., $f(r_{h})=0$. The mass $M$ of the black brane can be determined by applying this condition:
$m_{0}=\frac{\kappa q_{1}Q^{2}}{2r_{h}}+\frac{r_{h}^{3}}{2L^{2}}.$ (22)
By plugging Eq.(22) in Eq.(20) we have:
$f_{0}(r)=\frac{r^{2}}{L^{2}}(1-\frac{r_{h}^{3}}{r^{3}})+\frac{\kappa q_{1}Q^{2}}{4r}(\frac{1}{r}-\frac{1}{r_{h}}).$ (23)
We impose the condition that Eq. (A) and Eq. (A) must be equivalent up to the first order of $q_{2}$. Consequently, $H_{1}$ is calculated as follows:
$H_{1}(r)=-12\mathit{k}f_{0}^{\prime\prime}(r)h_{0}^{\prime}(r)^{2}-32\mathit{k}f_{0}^{\prime}(r)h_{0}^{\prime}(r)h_{0}^{\prime\prime}(r)-8\mathit{k}rf_{0}^{\prime\prime}(r)h_{0}^{\prime}(r)h_{0}^{\prime\prime}(r)$
$+\frac{24\mathit{k}f_{0}^{\prime}(r)h_{0}^{\prime}(r)^{2}}{r}-4\mathit{k}rf_{0}^{(3)}(r)h_{0}^{\prime}(r)^{2}-\frac{16\mathit{k}f_{0}(r)h_{0}^{\prime}(r)h_{0}^{\prime\prime}(r)}{r}$
$+\frac{24\mathit{k}f_{0}(r)h_{0}^{\prime}(r)^{2}}{r^{2}}.$ (24)
Hence, $H_{1}$ is calculated as follows:
$H_{1}(r)=-\frac{80\kappa Q^{2}\Lambda}{r^{4}}=\frac{240\kappa Q^{2}}{L^{2}r^{4}}.$ (25)
$h_{1}(r)$ can be calculated as follows,
$h_{1}(r)$ $=C_{1}\int^{r}_{r_{h}}\left(\frac{H_{1}(u)}{q_{1}u^{2}}-\frac{4\left(u\left(4f_{0}^{\prime}(u)+uf_{0}^{\prime\prime}(u)\right)+2f_{0}(u)\right)^{2}}{q_{1}^{2}u^{6}}\right)du$ (26)
$=\frac{576C_{1}}{L^{4}q_{1}^{2}}(\frac{1}{r}-\frac{1}{r_{h}})-\frac{24\mathit{k}Q^{2}C_{1}}{L^{2}q_{1}}(\frac{1}{r^{5}}-\frac{1}{r_{h}^{5}}),$
where $C_{1}=q_{1}Q$.
$h_{1}(r)=\frac{576Q}{L^{4}q_{1}}(\frac{1}{r}-\frac{1}{r_{h}})-\frac{48\mathit{k}Q^{3}}{L^{2}}(\frac{1}{r^{5}}-\frac{1}{r_{h}^{5}}),$ (27)
$f_{1}(r)$ can be calculated as:
$f_{1}(r)=-\frac{672\mathit{k}m_{0}Q^{2}}{L^{2}r^{5}}+\frac{288\mathit{k}^{2}Q^{4}q_{1}}{L^{2}r^{6}}+\frac{768\mathit{k}Q^{2}}{L^{4}r^{2}}-\frac{2m_{1}}{r}.$ (28)
We have obtained the solution up to the linear order of $q_{2}$:
$h(r)=Q(\frac{1}{r}-\frac{1}{r_{h}})+q_{2}\Big\{\frac{576Q}{L^{4}q_{1}}(\frac{1}{r}-\frac{1}{r_{h}})-\frac{48\mathit{k}Q^{3}}{L^{2}}(\frac{1}{r^{5}}-\frac{1}{r_{h}^{5}})\Big\},$ (29)
$H(r)=-q_{2}\frac{80\kappa Q^{2}\Lambda}{r^{4}},$ (30)
$f(r)=-\frac{2m_{0}}{r}+\frac{r^{2}}{L^{2}}+\frac{\kappa q_{1}Q^{2}}{r^{2}}$
$+q_{2}\Big(-\frac{672\mathit{k}m_{0}Q^{2}}{L^{2}r^{5}}+\frac{288\mathit{k}^{2}Q^{4}q_{1}}{L^{2}r^{6}}+\frac{768\mathit{k}Q^{2}}{L^{4}r^{2}}-\frac{2m_{1}}{r}\Big).$ (31)
We reparametrize such that:
$m_{0}+q_{2}m_{1}=:\bar{m}_{0},$ (32)
So $f(r)$ changes as:
$f(r)=-\frac{2\bar{m}_{0}}{r}+\frac{r^{2}}{L^{2}}+\frac{\kappa q_{1}Q^{2}}{r^{2}}$
$+q_{2}\Big(-\frac{672\mathit{k}m_{0}Q^{2}}{L^{2}r^{5}}+\frac{288\mathit{k}^{2}Q^{4}q_{1}}{L^{2}r^{6}}+\frac{768\mathit{k}Q^{2}}{L^{4}r^{2}}\Big).$ (33)
This is the solution up to the linear order of $q_{2}$.
3 Holographic Complexity of $R^{2}F^{2}$
$Complexity=anything$ conjecture has emerged as a generalized version of the complexity in the bulk. It states that the complexity can be described by a more general functional to include a class of new diffeomorphism invariant observables [15]:
$\mathcal{O}_{F_{1},\Sigma_{F_{2}}}(\Sigma_{CFT})=\frac{1}{G_{N}L}\int_{\Sigma_{F_{2}}}\mathrm{d}^{d}\sigma\sqrt{h}F_{1}(g_{\mu\nu};X^{\mu})$ (34)
where, $F_{1}$ can be a general scalar function of metric $g_{\mu\nu}$ and an embedding $X^{\mu}(\sigma^{a})$ of the hypersurfaces. Moreover, $\Sigma_{F_{2}}$ could be a general hypersurface but we restrict ourself to codimension-one hypersurfaces in the bulk spacetime with boundary time slice $\partial\Sigma_{F_{2}}=\Sigma_{CFT}$. By extremizing the hypersurface it leads to:
$\delta_{X}\Big(\int_{\Sigma}\mathrm{d}^{d}\sigma\sqrt{h}F_{2}(g_{\mu\nu};X^{\mu})\Big)=0.$ (35)
For simplicity, we follow the case $F_{1}=F_{2}$, so the observable (34), known by $\mathcal{C}_{Any}$, obeying the above condition is given by:
$\mathcal{C}_{Any}(\tau)=\max_{\partial\Sigma(\tau)=\Sigma_{\tau}}\frac{V_{x}}{G_{N}L}\left[\int_{\Sigma}\mathrm{d}^{d}\sigma\sqrt{h}F_{1}(g_{\mu\nu};X^{\mu})\right]$ (36)
where $h$ is the determinant of induced metric on the given hypersurface. The usual choice of the generalization is:
$F_{1}=F_{2}=1+\gamma b,$ (37)
where $b$ could be $C^{2}$ the Weyl tensor squared, $F^{2}$, the field strength and $R^{2}F^{2}$, where $R$ is the Ricci scalar. For $\gamma=0$ the CV is retrieved. In the following, we are going to study the $\mathcal{C}_{Any}$ of this setup, with the metric solution (12). To do this, we need to transform the metric to Eddington-Finkelstein coordinates:
$\mathrm{d}s^{2}=-e^{-2H}f(r)\mathrm{d}v^{2}+2e^{-H}\mathrm{~d}v\mathrm{~d}r+\frac{r^{2}}{L^{2}}\big(dx^{2}+dy^{2}\big),$ (38)
where:
$v=t+r^{*}(r),\quad\mathrm{d}r^{*}=\frac{\mathrm{d}r}{f(r)e^{-H}}.$ (39)
Then, the complexity could be computed from (36) and (37) as [15]:
$\mathcal{C}_{Any}(\tau)=\frac{V_{0}}{G_{N}L}\int_{\Sigma}\mathrm{d}\sigma r^{d-1}\sqrt{-e^{-2H}f\dot{v}^{2}+2e^{-H}\dot{v}\dot{r}}a(r),$ (40)
where the dots indicate derivatives with respect to $\sigma$, and $V_{0}$ represents the volume of spatial $x$ and $y$ directions.
Supposing $\mathcal{C}_{Any}$ as an action, then the conserved momentum conjugate to coordinate $v$ would be:
$P_{v}=-\frac{\partial\mathcal{L}}{\partial\dot{v}}=\frac{a(r)r^{d-1}(-e^{-2H}f\dot{v}+e^{-H}\dot{r})}{\sqrt{-e^{-2H}f\dot{v}^{2}+2e^{-H}\dot{v}\dot{r}}}=\dot{r}-e^{-H}f(r)\dot{v}.$ (41)
Due to the fact that (40) is diffeomorphism invariant and doesn’t change under reparametrization, the parameter $\sigma$ could be fixed by choosing:
$\sqrt{-e^{-2H}f\dot{v}^{2}+2e^{-H}\dot{v}\dot{r}}=a(r)r^{d-1}e^{-H}.$ (42)
Then, extremality conditions is given:
$\dot{r}=\pm\sqrt{P^{2}_{v}+e^{-2H}f(r)a(r)^{2}r^{2(d-1)}},$ (43)
$\dot{t}=\dot{v}-\frac{\dot{r}}{e^{-H}f(r)}=\frac{-P_{v}\dot{r}}{e^{-H}f(r)\sqrt{P^{2}_{v}-\tilde{U}(r)}}.$ (44)
where $\tilde{U}(r)$ is an effective potential:
$\tilde{U}(r)=-e^{-2H}f(r)a(r)^{2}r^{2(d-1)}.$ (45)
Then, (43) gives the equation of motion like for a classical particle:
$\dot{r}^{2}+\tilde{U}(r)=P^{2}_{v}.$ (46)
On the symmetric trajectory, the conserved momentum is a function of the turning point $r_{min}$:
$P^{2}_{v}=\tilde{U}(r_{min})=-e^{-2H(r_{min})}f(r_{min})a(r_{min})^{2}r_{min}^{2(d-1)}.$ (47)
The boundary time is related to the conserved momentum and the turning point $r_{min}$ acts as an intermediary:
$\tau$ $=2\int_{r_{\min}}^{\infty}\mathrm{d}r\frac{-P_{v}}{e^{-H}f(r)\sqrt{P^{2}_{v}-\tilde{U}(r)}}.$ (48)
Boundary time derivation of $\mathcal{C}_{Any}$ leads to:
$\frac{\mathrm{d}\mathcal{C}_{Any}}{\mathrm{~d}\tau}=\frac{1}{2}\frac{\mathrm{d}\mathcal{C}_{Any}}{\mathrm{~d}\tau_{\mathrm{R}}}$ $=\frac{V_{0}}{G_{N}L}P_{v}$
$=\frac{V_{0}}{G_{N}L}\sqrt{-e^{-2H(r_{min})}f(r_{min})}a(r_{min})r_{min}^{(d-1)}.$ (49)
The growth rate of the generalized complexity can be considered by the behavior of the conserved momentum in full-time. The late-time behavior of the growth rate is given by:
$\lim_{\tau\rightarrow\infty}\frac{\mathrm{d}\mathcal{C}_{Any}}{\mathrm{~d}\tau}=\frac{V_{0}}{G_{N}L}\sqrt{-e^{-2H(\tilde{r}_{min})}f\left(\tilde{r}_{\min}\right)}a(\tilde{r}_{\min})\tilde{r}_{min}^{(d-1)},$ (50)
where $\tilde{r}_{\min}$ is local maximum of the effective potential, and owing to the fact that at this radius time goes to infinity, it is known as $r_{f}:=\tilde{r}_{\min}$.
The numerical analysis of the $\mathcal{C}_{Any}$-boundary time diagrams for the Weyl squared tensor $a^{2}=C^{2}$ are plotted in Fig.(1) and the complexity-boundary time in Fig.(2), for $R^{2}F^{2}$ are Fig.s (7), (4) and for $F^{2}$ in Fig.s(5), (6). Regarding the fact that the complexity being defined as a maximum value, the intersections of the complexity-boundary time diagrams indicate points where complexity switches from one upper branch to another, known as $\tau_{turning}$. At this time, a transition occurs in $\mathcal{C}_{Any}$-boundary time diagrams.
Figure 1: CGR in terms of the boundary time for $b=C^{2}$ (a) with the parameters Q=0.09Q=-0.09, $q_{2}=-0.06$ and $\gamma=-0.002$. (b) Q=0.09Q=-0.09, $q_{2}=-0.03$ and $\gamma=-0.002$. (c) Q=0.09Q=-0.09, $q_{2}=-0.03$ and $\gamma=0.002$. (d) Q=0.1Q=-0.1, $q_{2}=-0.03$ and $\gamma=-0.002$.
Figure 2: Complexity in terms of the boundary time for $b=C^{2}$ (a) with the parameters Q=0.09Q=-0.09, $q_{2}=-0.06$ and $\gamma=-0.002$. (b) Q=0.09Q=-0.09, $q_{2}=-0.03$ and $\gamma=-0.002$. (c) Q=0.09Q=-0.09, $q_{2}=-0.03$ and $\gamma=0.002$. (d) Q=0.1Q=-0.1, $q_{2}=-0.03$ and $\gamma=-0.002$.
Figure 3: CGR in terms of the boundary time for $b=R^{2}F^{2}$ (a) with the parameters Q=0.09Q=-0.09, $q_{2}=-0.06$ and $\gamma=-0.002$. (b) Q=0.09Q=-0.09, $q_{2}=-0.03$ and $\gamma=-0.002$. (c) Q=0.09Q=-0.09, $q_{2}=-0.03$ and $\gamma=0.002$. (d) Q=0.1Q=-0.1, $q_{2}=-0.03$ and $\gamma=-0.002$.
Figure 4: Complexity in terms of the boundary time for $b=R^{2}F^{2}$ (a) with the parameters Q=0.09Q=-0.09, $q_{2}=-0.06$ and $\gamma=-0.002$. (b) Q=0.09Q=-0.09, $q_{2}=-0.03$ and $\gamma=-0.002$. (c) Q=0.09Q=-0.09, $q_{2}=-0.03$ and $\gamma=0.002$. (d) Q=0.1Q=-0.1, $q_{2}=-0.03$ and $\gamma=-0.002$.
Figure 5: CGR in terms of the boundary time for $b=F^{2}$ (a) with the parameters Q=0.09Q=-0.09, $q_{2}=-0.06$ and $\gamma=-0.002$. (b) Q=0.09Q=-0.09, $q_{2}=-0.03$ and $\gamma=-0.002$. (c) Q=0.09Q=-0.09, $q_{2}=-0.03$ and $\gamma=0.002$. (d) Q=0.1Q=-0.1, $q_{2}=-0.03$ and $\gamma=-0.002$.
Figure 6: Complexity in terms of the boundary time for $b=F^{2}$ (a) with the parameters Q=0.09Q=-0.09, $q_{2}=-0.06$ and $\gamma=-0.002$. (b) Q=0.09Q=-0.09, $q_{2}=-0.03$ and $\gamma=-0.002$. (c) Q=0.09Q=-0.09, $q_{2}=-0.03$ and $\gamma=0.002$. (d) Q=0.1Q=-0.1, $q_{2}=-0.03$ and $\gamma=-0.002$.
4 Physical interpretation of CGR
As Fig.s(2), (4) and (6) show, by adopting different parameters of $Q$, $q_{2}$ and $\gamma$, the diagrams differ in the locations of the branches and the distance between the branches. The main question is that if the differences in the diagrams of CGR have a physical origin or not. In [38], it shows that the locations of the late time CGR (where the branches are divided) are controlled by the value of the bulk fields in the case $b=C^{2}$. For example, if there is an scalar field in the bulk, at the locations of the late time CGR, $r_{i}$, there is the following relation [38]:
$\Big(\nabla_{i}<T_{jk}>\Big)^{2}\sim\partial_{i}\phi^{(0)}\partial^{i}\phi^{(0)}=\frac{1-g^{(0)}_{tt}}{2\gamma g^{(0)}_{tt}}$, (51)
where $T_{jk}$ is the energy-momentum tensor at $r_{i}$, the location of the late time of CGR, $\phi^{(0)}$ is the value of the scalar field at the cut-off $r_{i}$ and $g_{tt}^{(0)}$ is the $tt$-component of the metric (12) in Fefferman-Graham coordinates at the cut-off $r_{i}$.
The same reasoning holds for the other choices of $b$.
4.1 The generalization term $\gamma b$
In this section the generalization term is interpreted as the penalty factor in the boundary circuit.
Let us rewrite (40) as follows:
$\mathcal{C}_{Any}\sim\int d\sigma r^{d-1}a\sqrt{g_{ij}\frac{\partial x^{i}}{\partial\sigma}\frac{\partial x^{j}}{\partial\sigma}}=\int d\sigma\sqrt{(r^{2(d-1)}a^{2}g_{ij})\dot{x}^{i}\dot{x}^{j}}=\int d\sigma\sqrt{g^{\prime}_{ij}\dot{x}^{i}\dot{x}^{j}}$ (52)
The above relation is similar to Nielson geometrical method to compute complexity in the quantum theory [39]:
$\mathcal{C}_{QFT}=\int dt\sqrt{\sum_{\alpha}\gamma_{\alpha}(Y^{\alpha})^{2}}=\int ds\sqrt{G_{IJ}\dot{Y}^{I}\dot{Y}^{J}}$ (53)
where the metric is defined in a Hilbert space of the unitary operators:
$ds^{2}=\sum_{\alpha}\gamma_{\alpha}(Y^{\alpha})^{2}dt^{2}.$ (54)
To make the cost function, $G_{IJ}$ in (53) acting differently in directions, there introduces a penalty factor which weighting some gates heavily and some lightly by a coefficient into the cost function. In the way of mere similarity, it seems that the coefficient $(r^{d-1}a)^{2}$ in (52) would be a position-dependent penalty factor or let us call it curvature penalty factor.
The mere similarity could be broaden by the bulk symplectic form and Berry curvature form in gauge/gravity duality. Let us starting with $\mathcal{C}_{QFT}$ and the duality of the variation of gravitational observables (which is the bulk symplectic form) and the Berry curvature form [40]. The Berry curvature form could be written in terms of the imaginary part of the quantum geometric tensor or Fubini-Study (FS) metric [41] in the boundary theory :
$\delta\mathcal{C}_{Any}=\Omega=\omega=\Im\mathcal{G}_{IJ},$ (55)
$\Omega$ is the bulk symplectic form and $\omega$ the Berry curvature form. Furthermore, FS metric is related to $G_{IJ}$ as:
$G_{IJ}=\mathcal{P}_{IM}\mathcal{G}^{MN}\mathcal{P}_{NJ},$ (56)
where $\mathcal{P}_{IJ}$ is the penalty factor matrix. One could rewrite (53) in terms of the bulk symplectic form as follows:
$\int ds\sqrt{\mathcal{P}_{IM}\mathcal{G}^{MN}\mathcal{P}_{NJ}\dot{Y}^{I}\dot{Y}^{J}}$ $=\int ds\sqrt{\mathcal{P}_{IM}(\Re{\mathcal{G}^{MN}}+i\Im{\mathcal{G}^{MN}})\mathcal{P}_{NJ}\dot{Y}^{I}\dot{Y}^{J}},$ (57)
$=\int ds\sqrt{\mathcal{P}_{IM}(\Re{\mathcal{G}^{MN}}+i\delta\mathcal{C}_{Any})\mathcal{P}_{NJ}\dot{Y}^{I}\dot{Y}^{J}},$ (58)
$=\int ds\sqrt{\mathcal{P}_{IM}(\Re{\mathcal{G}^{MN}}+i\delta\int d\sigma\sqrt{g^{\prime}_{ij}\dot{x}^{i}\dot{x}^{j}})\mathcal{P}_{NJ}\dot{Y}^{I}\dot{Y}^{J}}.$ (59)
Then,
$\mathcal{C}_{QFT}=\int ds\sqrt{\mathcal{P}_{IM}\mathcal{P}_{NJ}\Re{\mathcal{G}^{MN}}\dot{Y}^{I}\dot{Y}^{J}+i\mathcal{P}_{IM}\mathcal{P}_{NJ}\delta\int d\sigma\sqrt{g^{\prime}_{ij}\dot{x}^{i}\dot{x}^{j}}\dot{Y}^{I}\dot{Y}^{J}}.$ (60)
In the last line the factor $\mathcal{P}_{IM}\mathcal{P}_{NJ}\delta\int d\sigma\sqrt{g^{\prime}_{ij}\dot{x}^{i}\dot{x}^{j}}$ could be seen as a source for the penalty factor $\mathcal{P}(a)$, where $a$ is the generalization factor in (52). By change of $a$ the penalty factor in $\mathcal{C}_{QFT}$ is deformed.
As a result, by the change of $\gamma b$, $\mathcal{C}_{Any}$ and $\dot{\mathcal{C}}_{Any}$ are changed. A question may arise: is there a penalty factor in quantum circuits that its change makes the counterpart of $\mathcal{C}_{Any}$ as the gate complexity and $\dot{\mathcal{C}}_{Any}$ as the depth of the circuit [9] changed? we introduce two of these circuits in the section 4.4 which are applicable in quantum computations.
4.2 The non-minimal coupling $q_{2}$
In this section we argue that the non-minimal coupling $q_{2}$ is addressed to the scrambling time of the boundary theory. The non-minimal term acts as a deformation of Einstein-Maxwell theory. The dual of this theory in two dimensions (by dimension reduction and some compactifications) is Sachdev-Ye-Kitaev (SYK) model [^1] We should note that the duality of SYK model and black holes (Jackiew-Teitelboim gravity) is well-established for the $q_{2}=0$, without the non-minimal coupling. With $q_{2}\neq 0$ the dual theory would be a deformed SYK model; nonetheless, at low-curvature regime it reduces to SYK model. In addition, our results about the curvature penalty factor is intact whether $q_{2}=0$ or not. which has a very little scrambling time. By deformation, we are going away from the fast scrambler but this not a simple relationship.
The scrambling time of the bulk theory is independent of the generalization parameter. However, it seems that the scrambling time should depend on the penalty factor in the quantum circuit, because of the fact that the penalty factor in the circuit makes an effect on the scrambling of information. Nonetheless, in the bulk the role of the penalty factor is ignored. In the following we define an effective scrambling time in the bulk theory call it the new scrambling time and show that how it controls CGR.
As (60) shows the action of the particle in (40) affects in the complexity of the boundary theory. However, by the generalized term the metric is scaled [^2] It is worth mentioning that this is a scaling and not the transformation of the coordinates.:
$ds^{2}_{scal.}:=-a(r)^{2}g_{tt}dt^{2}+a(r)^{2}g_{rr}dr^{2}+r^{2}(dx^{2}+dy^{2}),$ (61)
where $g_{xx}$ is not scaled. In this way, it seems that the scaled metric or the generalized term weights different directions unequally. But we should have it in mind that the rescaling does not change the bulk theory and only the definition of the complexity. In other words, the butterfly velocity and the scrambling time would be intact. In this section we introduce a new butterfly velocity and the scrambling time which is due to the action (40) or (61), where the dynamic of the particle determines the complexity. Put it differently, when the metric is scaled by the generalized term in the complexity functional, the new scrambling time would be changed which signaling the unequal weight of the particle motion.
The new butterfly velocity in $d+1$ dimensions in (61) is determined by [42]:
$v_{(scal.)B}^{2}=\frac{(a^{2}g_{tt})^{\prime}}{(d-1)g^{\prime}_{xx}},$ (62)
where the derivatives are with respect to $r$. Then, the new butterfly velocity $v_{(scal.)B}$ is scaled as:
$v_{(scal.)B}^{2}$ $=\frac{1}{a^{2}}\Big|_{r=R=1}v_{B}^{2}.$ (63)
Besides the scrambling time (B-15) for the action (40), we introduce the new scrambling time which is scaled as:
$t_{scal.}\approx\frac{x}{v_{(scal.)B}}=\frac{t_{*}}{|a(R)|}.$ (64)
It is believed that there is an inverse relation between the scrambling time and the complexity growth rate, it means that by an increase in the scrambling time, the processing rate would be decreased. However, there is not a universal relation, it is model-dependent. The leading term in (64) would be:
$t_{scal.}\approx\frac{1}{\sqrt{g^{\prime}_{0}(R)}}\Big(1-\frac{q_{2}}{2}\frac{g^{\prime}_{1}(R)}{g^{\prime}_{0}(R)}-\gamma b_{0}(R)-\gamma q_{2}b_{1}(R)+\frac{\gamma q_{2}b_{0}(R)}{2}\frac{g^{\prime}_{1}(R)}{g^{\prime}_{0}(R)}\Big),$ (65)
where $g_{tt}=g_{0}+q_{2}g_{1}$ and $b=b_{0}+q_{2}b_{1}$. In addition, the leading term of the complexity growth rate in (49) would be:
$\dot{\mathcal{C}}_{Any}(r_{f})$ $\sim a(r_{f})\sqrt{g_{tt}(r_{f})},$ (66)
$\approx\sqrt{g_{0}(r_{f})}\Big(1+\frac{q_{2}}{2}\frac{g_{1}(r_{f})}{g_{0}(r_{f})}+\gamma b_{0}(r_{f})+\gamma q_{2}b_{1}(r_{f})+\frac{\gamma q_{2}b_{0}(r_{f})}{2}\frac{g_{1}(r_{f})}{g^{\prime}_{0}(r_{f})}\Big).$ (67)
By the following approximations near the horizon $r_{f}\approx R$:
$g^{\prime}_{tt}(R)(r_{f}-R)\approx g_{tt}(r_{f}),$ (68)
$g^{\prime}_{0}(R)(r_{f}-R)\approx g_{0}(r_{f}),$ (69)
$g^{\prime}_{1}(R)(r_{f}-R)\approx g_{1}(r_{f}).$ (70)
(67) is given:
$\frac{\sqrt{r_{f}-R}}{\dot{\mathcal{C}}_{Any}(r_{f})}\approx\frac{1-\frac{q_{2}}{2}\frac{g^{\prime}_{1}(R)}{g^{\prime}_{0}(R)}-\gamma b_{0}(r_{f})-\gamma q_{2}b_{1}(r_{f})+\frac{\gamma q_{2}b_{0}(r_{f})}{2}\frac{g^{\prime}_{1}(R)}{g^{\prime}_{0}(R)}}{\sqrt{g^{\prime}_{0}(R)}}.$ (71)
Putting (71) into (65), we could write the new scrambling time in terms of the complexity growth rate:
$t_{scal.}$ $\approx\frac{\sqrt{r_{f}-R}}{\dot{\mathcal{C}}_{Any}(r_{f})}+\frac{\gamma}{\sqrt{g^{\prime}_{0}(R)}}\Big(b^{\prime}_{0}(R)+q_{2}b^{\prime}_{1}(R)\Big)(r_{f}-R)$
$+\frac{\gamma q_{2}}{\sqrt{g^{\prime}_{0}(R)}}\frac{g^{\prime}_{1}(R)}{g^{\prime}_{0}(R)}\Big(b_{0}(R)\Big),$ (72)
where $b^{\prime}_{0}(R)(r_{f}-R)\approx b_{0}(r_{f})-b_{0}(R)$. Then, the complexity growth rate near the horizon could be rewritten in terms of the bulk scrambling time:
$\dot{\mathcal{C}}_{Any}(r_{f})$ $\approx\frac{\sqrt{r_{f}-R}}{\frac{t_{*}}{|a(R)|}-\frac{\gamma}{\sqrt{g^{\prime}_{0}(R)}}\Big(b^{\prime}(R)\Big)(r_{f}-R)-\frac{\gamma q_{2}}{\sqrt{g^{\prime}_{0}(R)}}\frac{g^{\prime}_{1}(R)}{g^{\prime}_{0}(R)}\Big(b_{0}(R)\Big)},$
$\approx\frac{|1+\gamma b(R)|\sqrt{r_{f}-R}}{t_{*}-\frac{\gamma}{\sqrt{g^{\prime}_{0}(R)}}\Big(b^{\prime}(R)\Big)(r_{f}-R)-\frac{\gamma q_{2}}{\sqrt{g^{\prime}_{0}(R)}}\frac{g^{\prime}_{1}(R)}{g^{\prime}_{0}(R)}\Big(b_{0}(R)\Big)},$ (73)
$\approx\frac{1}{t_{scal.}}=\frac{|a(R)|}{t_{*}}$ (74)
The case $b=C^{2}$
In this case we have:
$b_{0}(R)=12-24Q^{2}+12Q^{4},$ (75)
$b^{\prime}_{0}(R)=-72+192Q^{2}-120Q^{4},$ (76)
$g^{\prime}_{0}(R)=3-Q^{2},$ (77)
$g^{\prime}_{1}(R)=1056Q^{2}.$ (78)
By (73) and setting $\gamma=-0.002$, we have:
$\dot{\mathcal{C}}_{Any}$ $\approx\frac{1+\gamma q_{2}}{(1+0.004q_{2})(-0.0832509-4.6865q_{2}+0.0396679q_{2})},$
$\approx 1+q_{2}\approx 1-|q_{2}|.$ (79)
Then, by the increase in $|q_{2}|$, as a result, there would be an increase in the new scrambling time, the complexity growth rate is decreased as Fig.(1a) and (1b) demonstrate.
The case $b=R^{2}F^{2}$
$b_{0}(R)$
, $b^{\prime}_{0}(R)$ and $b_{1}(R)$ for $b=R^{2}F^{2}$ would be:
$b_{0}(R)=-288Q^{2},$ (80)
$b^{\prime}_{0}(R)=1952Q^{2},$ (81)
$b_{1}(R)=-6144(54Q^{2}-84Q^{4}+21Q^{6})q_{2}.$ (82)
The complexity growth rate for $b=R^{2}F^{2}$ according to (73), (80), (82) and the fact that $\gamma=-0.002$ is given:
$\dot{\mathcal{C}}_{Any}$ $\approx\frac{1+\gamma(-288Q^{2}-2653.59q_{2})}{1+0.004q_{2}-\Big(0.3312-12.0417q2\Big)},$ (83)
$\approx 1-q_{2}=1+|q_{2}|$ (84)
As Fig.s(3a) and (3b) for Q=0.09Q=-0.09 show the complexity growth rate is decreased by the increase of $|q_{2}|$.
The case $b=F^{2}$
The leading term in $b_{0}(R)$, $b^{\prime}_{0}(R)$ and $b_{1}(R)$ for $b=F^{2}$ would be:
$b_{0}(R)=-2Q^{2},$ (85)
$b^{\prime}_{0}(R)=8Q^{2},$ (86)
$b_{1}(R)=1280Q^{4}-2304Q^{2}.$ (87)
The leading order in CGR based on (73), (85), (86) and setting $\gamma=-0.002$ would be:
$\dot{\mathcal{C}}_{Any}$ $\approx\frac{1+\gamma(-2Q^{2}-2304q_{2})}{1+0.004q_{2}-\gamma\Big(0.0374629-10.7893q2\Big)},$ (88)
$\approx 1+q_{2}=1-|q_{2}|$ (89)
Fig.s (5b) and (5d) illustrate that CGR is reduced by the increase in the non-minimal coupling.
In this section we showed that the new scrambling time is related to the non-minimal coupling $q_{2}$ and consequently, by the change of the coupling, CGR is changed as is expected.
4.3 The conserved charge $Q$
In this section we argue that the value of the conserved charge $Q$ restricts the simple unitary operators on the boundary, as result, it restricts the complexity and CGR.
[…]
  1. 1.
    $[F,G]=iH$,
  2. 2.
[…]
This lemma brings a restriction on the choice of the simple unitary operators. Due to the fact that the accuracy $\epsilon$ should lie in the interval $\Big(0,max[2\sin(|E|/2)]\Big]$, it demonstrates that the eigenvalue of the operator $U$ could affect the choice of the quantum gates.
In this manner, the distance for a particular accuracy is restricted by the value of the conserved charge $Q$, which means that the change in the value of the charge restricts the simple unitary operators and as consequence CGR is restricted.
However, it could be shown approximately that the new scrambling time for the solution (12) has an approximate relation with the conserved charge. Then, by the increase of the value of the conserved charge, the new scrambling time would be increased and as a result the rate of the complexity as the rate of the processing would be decreased.
The case $b=C^{2}$
For $b=C^{2}$: The complexity growth rate based on (73), (75) and (76) is given by:
$\dot{\mathcal{C}}_{Any}(r_{f})$ $\approx\frac{|1+\gamma 12(1-2Q^{2}+Q^{4})|\sqrt{r_{f}-R}}{t_{*}-\frac{2\gamma}{\sqrt{3}}(-72+180Q^{2})(r_{f}-R)-\frac{\gamma q_{2}(1056Q^{2})}{(3-Q^{2})^{\frac{3}{2}}}(12-24Q^{2})},$ (90)
$\approx\frac{1-2\gamma Q^{2}}{1+\frac{Q^{2}}{2}}\approx 1-\frac{Q^{2}}{2}.$ (91)
Then, by increase in the conserved charge we see a decrease in the complexity growth rate.
The case $b=R^{2}F^{2}$
The complexity growth rate (73) would be:
$\dot{\mathcal{C}}_{Any}(r_{f})$ $\approx\frac{1-288\gamma Q^{2}}{1+\frac{Q^{2}}{2}}\approx 1-\frac{Q^{2}}{2}$ (92)
Fig.s (3b) and (3d) displays that by increase in the conserved charge, CGR is reduced.
The case $b=F^{2}$
CGR according to (73) is given by:
$\dot{\mathcal{C}}_{Any}(r_{f})$ $\approx\frac{1-2304\gamma Q^{2}}{1+\frac{Q^{2}}{2}}\approx 1-\frac{Q^{2}}{2}$ (93)
Fig.s (5b) and (5d) depict the decrease of CGR by the increase in the conserved charge.
All in all, CGR is reduced by the increase of the conserved charge as is expected of the limitation of the quantum gates in SKA.
4.4 Superconducting qubits circuit
In this section, we provide two examples in superconducting qubits circuits where by an increase in the penalty factor, CGR would be increased or decreased, which is dependent to the penalty factor. The first example shows a direct relation between the penalty factor and CGR: a quantum circuit that CGR is increased by the increasing the penalizing the circuit. Moreover, the second one provides an inverse relation.
In the superconducting qubit circuits to prepare a qubit one needs controlling quickly the fast gates (or low runtime) to stay away from decoherence. However, fast controlling pose a problem because of the fact that it requires a strong, sharp microwave pulse, but strong pulse has a wide frequency spectrum. A strong pulse allows for very fast $CNOT$ gates [45]. Nonetheless, this makes parallelizing operations to reduce circuit depth. Moreover, a wide frequency pulse excite the qubit to higher energy states which leads to the leakage error or crosstalk between energy levels. To reduce the leakage it requires the qubit be more anharmonic that means the energy gap of the states be more significantly different [45]. The anharmonicity makes the states more isolated from the states with higher energies. A problem back of the other, the anharmonicity makes the qubit more sensitive to charge noise which results in reducing the coherence time. As a way out, one needs increase the circuit depth by applying crosstalk penalty.
Increase in crosstalk penalty, $\gamma_{\alpha}$ means that the algorithm weights heavily the parallel operations which leads to the crosstalk and makes it sequentially instead, as a consequence, increase in the circuit depth, $\dot{\mathcal{C}}_{circ.}$. To avoid the possible errors, insertion of correction pulses lets to mitigate the effects of errors which increases the gate count, $\mathcal{C}_{circ.}$. In the following we address a concrete example of the crosstalk penalty.
Suppose that there are three qubits of QB1 connected to QB2 and QB2 connected to QB3 where crosstalk is high between QB1&\&QB2 and QB2&\&QB3. If one needs to apply a CNOT(1,3) operator, it is not possible due to the fact that they are not adjacent. Then, it should be applied a SWAP(2,3) operation and then applies CNOT(1,3). However, SWAP(2,3) is not a primitive gate. SWAP is complied into three CNOT operators as the figure (7) shows. Similar occasion is appeared in IBM superconducting quantum circuit [46].
Figure 7: A schematic quantum circuit with three qubits. The time goes from left to right and each horizontal line depicts a qubit. In addition, the dotted line containing a circle shows an operation or quantum gate and the simple vertical dotted line, the time step. (a) a circuit that applies a CNOT operation to qubit 1 and 3 in one step of time. (b) a circuit that operates a SWAP gate to qubit 2 and 3 in a time step and then applies a CNOT gate to qubit 1 and 3 in the next time step.
Another example that demonstrates an inverse relation between the penalizing and CGR is in the quantum approximation optimization algorithm (QAOA) where to optimize the algorithm inserts some constraints (as a penalty factor). These constraints make the algorithm more efficient and reduce the circuit depth as an optimization. These algorithm are used in superconducting circuits that are prone to noises and require shallower circuits to prevent noise decohrence. If the evolution of the problem is described by problem Hamiltonian or the system Hamiltonian, then the constraints as an optimization is added to system:
$H_{optimized}=H_{system}+\gamma_{a}H_{constraints},$ (94)
where $\gamma_{a}$ parameter tune the intense of the optimization or penalizing (similar to $\gamma$ in the generalized complexity). Then, the optimization would be the qubit that has the most overlap with the lower eigenstate of the system after $p$ layers of operations or depth $p$. [47] shows such a relation where the intensifying $\gamma_{a}$ makes $p$ layers reduced. It means that by intensifying the obeying rules, the system in shorter steps gets to the desired solution.
By the way, in [48], the authors provide an algorithm known as mixer-phaser ansatz in superconducting processors in which they did not used a penalty-based algorithm. In a layman’s term, if the problem system goes away from your desired solution, the penalty factor penalizes the system to obey the constraints. However, in [48] the constraints have been inserted before the system circuit. Then, they get rid of the penalizing.
The analogue of the superconducting circuits is SYK model that is a solvable model of strange metal states where it has a non-local Hamiltonian [49]. Now, instead of qubits, we are dealing with some strongly interacting fermions, then the Hilbert space of fermions in SYK model could be mapped to the Hilbert space of qubits by Bravyi-Kitaev transformation [50] and Jordan-Wigner transformation [51]. Put it differently, these examples could be transformed into strange metal circuits with fermions.
5 Conclusion
In this study, we presented a black brane solution in four-dimensional AdS spacetime Einstein-Maxwell theory which is characterized by a non-minimal term $R^{2}F_{\mu\nu}F^{\mu\nu}$. As this model lacks a non-perturbative solution, we addressed it using a perturbative approach based on the non-minimal coupling $q_{2}$ up to first order. This non-minimal coupling can partially describe strange metals; specifically, our model exhibits a linear relationship between the resistivity and the temperature, though it does not include a phase transition. As a consequence, this model could be used to implement quantum computations. Although, superconducting processors are widely used in quantum computations, there are some transformations that make a connection between the qubits of the superconducting circuits and the fermions of the strange metal circuits such as Bravyi-Kitaev and Jordan-Wigner transformations.
We considered the complexity of this model via the complexity equals anything proposal. In this proposal there is an arbitrariness of the generalization parameter $b$. The computations have been done for three choice parameters of $C^{2}$, the Weyl squared tensor, $R^{2}F^{2}$, the squared of the Ricci scalar and the Maxwell field strength, and $F^{2}$. CGR in this model is characterized by three parameters of the conserved charge $Q$, the non-minimal coupling $q_{2}$ and the generalized parameter $b$. We showed that:
  • The generalization parameter acts as a penalty factor in the bulk.
  • The non-minimal coupling is addressed to the effective scrambling time that contains the generalization parameter.
  • The conserved charge is related to Solovay-Kitaev algorithm that states the increase in the value of the conserved charge makes the available simple unitary operators restricted and as consequence the restriction of the processing rate or CGR.
It has been seen that the three generalization parameters have different behavior in change of the non-minimal coupling. CGR in $C^{2}$ and $F^{2}$ is decreased by the $q_{2}$ increase and $R^{2}F^{2}$ by $q_{2}$ diminishment. In other words, a direct and an inverse relationship. Then, we provided two quantum circuit examples in the superconducting quantum circuits that show the behavior of CGR in our solution. These examples are widely applied in superconducting-based computations and demonstrate how intensifying the penalty factor in the circuits could suppressed or encouraged CGR. Moreover, we have noticed that CGR with the generalization parameter $F^{2}$ does not contain branches and its behavior is similar to CV and CA.
In addition, in quantum algorithm has been introduced some algorithms that by removing the penalty factor received more efficient circuits with shorter depth circuits. For future research we think that this could inspire to introduce a holographic complexity proposal that is free of penalizing and restricts the target states instead of penalizing the simple unitary operators. Furthermore, it seems that the generalization parameter of the kind of spacetime quantities like $R$, $R^{2}$ and $C^{2}$ lead to different branches in CGR. For future research this point could be scrutinized that why there is no branch in CGR when the generalization parameter is independent of spacetime quantities.
Acknowledgment
Authors would like to express their gratitude to Monireh Emami for her valuable help in the numerical calculations.
Appendix A Equations of motion
The $tt$-component of the equations of motion is as follows:
$20q_{2}f(r)^{3}h^{\prime}(r)^{2}H^{\prime}(r)^{4}+28q_{2}f(r)^{2}f^{\prime}(r)h^{\prime}(r)^{2}H^{\prime}(r)^{3}+\frac{16q_{2}f(r)^{3}h^{\prime}(r)^{2}H^{\prime}(r)^{3}}{r}$
$+64q_{2}f(r)^{3}h^{\prime}(r)h^{\prime\prime}(r)H^{\prime}(r)^{3}-35q_{2}f(r)f^{\prime}(r)^{2}h^{\prime}(r)^{2}H^{\prime}(r)^{2}$
$-\frac{160q_{2}f(r)^{2}f^{\prime}(r)h^{\prime}(r)^{2}H^{\prime}(r)^{2}}{r}-\frac{32q_{2}f(r)^{3}h^{\prime}(r)^{2}H^{\prime}(r)^{2}}{r^{2}}$
$+16q_{2}f(r)^{3}h^{\prime\prime}(r)^{2}H^{\prime}(r)^{2}-36q_{2}f(r)^{2}h^{\prime}(r)^{2}f^{\prime\prime}(r)H^{\prime}(r)^{2}$
$-56q_{2}f(r)^{2}f^{\prime}(r)h^{\prime}(r)h^{\prime\prime}(r)H^{\prime}(r)^{2}-\frac{96q_{2}f(r)^{3}h^{\prime}(r)h^{\prime\prime}(r)H^{\prime}(r)^{2}}{r}$
$+72q_{2}f(r)^{3}h^{\prime}(r)^{2}H^{\prime\prime}(r)H^{\prime}(r)^{2}+16q_{2}f(r)^{3}h^{\prime}(r)h^{(3)}(r)H^{\prime}(r)^{2}$
$+\frac{68q_{2}f(r)f^{\prime}(r)^{2}h^{\prime}(r)^{2}H^{\prime}(r)}{r}+\frac{152q_{2}f(r)^{2}f^{\prime}(r)h^{\prime}(r)^{2}H^{\prime}(r)}{r^{2}}-24q_{2}f(r)^{2}f^{\prime}(r)h^{\prime\prime}(r)^{2}H^{\prime}(r)$
$-\frac{32q_{2}f(r)^{3}h^{\prime\prime}(r)^{2}H^{\prime}(r)}{r}+16q_{2}f(r)f^{\prime}(r)h^{\prime}(r)^{2}f^{\prime\prime}(r)H^{\prime}(r)+\frac{64q_{2}f(r)^{2}h^{\prime}(r)^{2}f^{\prime\prime}(r)H^{\prime}(r)}{r}$
$-12q_{2}f(r)f^{\prime}(r)^{2}h^{\prime}(r)h^{\prime\prime}(r)H^{\prime}(r)+\frac{64q_{2}f(r)^{3}h^{\prime}(r)h^{\prime\prime}(r)H^{\prime}(r)}{r^{2}}-16q_{2}f(r)^{2}h^{\prime}(r)f^{\prime\prime}(r)h^{\prime\prime}(r)H^{\prime}(r)$
$-108q_{2}f(r)^{2}f^{\prime}(r)h^{\prime}(r)^{2}H^{\prime\prime}(r)H^{\prime}(r)-\frac{144q_{2}f(r)^{3}h^{\prime}(r)^{2}H^{\prime\prime}(r)H^{\prime}(r)}{r}+4q_{2}f(r)^{2}h^{\prime}(r)^{2}f^{(3)}(r)H^{\prime}(r)$
$-24q_{2}f(r)^{2}f^{\prime}(r)h^{\prime}(r)h^{(3)}(r)H^{\prime}(r)-\frac{32q_{2}f(r)^{3}h^{\prime}(r)h^{(3)}(r)H^{\prime}(r)}{r}-16q_{2}f(r)^{3}h^{\prime}(r)^{2}H^{(3)}(r)H^{\prime}(r)$
$-\frac{36q_{2}f(r)f^{\prime}(r)^{2}h^{\prime}(r)^{2}}{r^{2}}+\frac{1}{4}q_{1}f(r)h^{\prime}(r)^{2}-\frac{48q_{2}f(r)^{2}f^{\prime}(r)h^{\prime}(r)^{2}}{r^{3}}+\frac{12q_{2}f(r)^{3}h^{\prime}(r)^{2}}{r^{4}}$
$-3q_{2}f(r)h^{\prime}(r)^{2}f^{\prime\prime}(r)^{2}+\frac{32q_{2}f(r)^{2}f^{\prime}(r)h^{\prime\prime}(r)^{2}}{r}+8q_{2}f(r)^{2}f^{\prime\prime}(r)h^{\prime\prime}(r)^{2}+\frac{16q_{2}f(r)^{3}h^{\prime\prime}(r)^{2}}{r^{2}}$
$-12q_{2}f(r)^{3}h^{\prime}(r)^{2}H^{\prime\prime}(r)^{2}+\frac{e^{-2H(r)}\Lambda f(r)}{\mathit{k}}+\frac{e^{-2H(r)}f(r)f^{\prime}(r)}{\mathit{k}r}-\frac{12q_{2}f(r)f^{\prime}(r)h^{\prime}(r)^{2}f^{\prime\prime}(r)}{r}$
$+\frac{16q_{2}f(r)f^{\prime}(r)^{2}h^{\prime}(r)h^{\prime\prime}(r)}{r}+\frac{40q_{2}f(r)^{2}f^{\prime}(r)h^{\prime}(r)h^{\prime\prime}(r)}{r^{2}}-\frac{32q_{2}f(r)^{3}h^{\prime}(r)h^{\prime\prime}(r)}{r^{3}}$
$+4q_{2}f(r)f^{\prime}(r)h^{\prime}(r)f^{\prime\prime}(r)h^{\prime\prime}(r)+\frac{80q_{2}f(r)^{2}h^{\prime}(r)f^{\prime\prime}(r)h^{\prime\prime}(r)}{r}-10q_{2}f(r)f^{\prime}(r)^{2}h^{\prime}(r)^{2}H^{\prime\prime}(r)$
$-\frac{8q_{2}f(r)^{2}f^{\prime}(r)h^{\prime}(r)^{2}H^{\prime\prime}(r)}{r}+\frac{32q_{2}f(r)^{3}h^{\prime}(r)^{2}H^{\prime\prime}(r)}{r^{2}}-16q_{2}f(r)^{3}h^{\prime\prime}(r)^{2}H^{\prime\prime}(r)$
$-12q_{2}f(r)^{2}h^{\prime}(r)^{2}f^{\prime\prime}(r)H^{\prime\prime}(r)-88q_{2}f(r)^{2}f^{\prime}(r)h^{\prime}(r)h^{\prime\prime}(r)H^{\prime\prime}(r)-\frac{96q_{2}f(r)^{3}h^{\prime}(r)h^{\prime\prime}(r)H^{\prime\prime}(r)}{r}$
$+2q_{2}f(r)f^{\prime}(r)h^{\prime}(r)^{2}f^{(3)}(r)+\frac{24q_{2}f(r)^{2}h^{\prime}(r)^{2}f^{(3)}(r)}{r}+16q_{2}f(r)^{2}h^{\prime}(r)h^{\prime\prime}(r)f^{(3)}(r)$
$+\frac{32q_{2}f(r)^{2}f^{\prime}(r)h^{\prime}(r)h^{(3)}(r)}{r}+\frac{16q_{2}f(r)^{3}h^{\prime}(r)h^{(3)}(r)}{r^{2}}+8q_{2}f(r)^{2}h^{\prime}(r)f^{\prime\prime}(r)h^{(3)}(r)$
$-16q_{2}f(r)^{3}h^{\prime}(r)H^{\prime\prime}(r)h^{(3)}(r)-32q_{2}f(r)^{2}f^{\prime}(r)h^{\prime}(r)^{2}H^{(3)}(r)-\frac{32q_{2}f(r)^{3}h^{\prime}(r)^{2}H^{(3)}(r)}{r}$
$-32q_{2}f(r)^{3}h^{\prime}(r)h^{\prime\prime}(r)H^{(3)}(r)+4q_{2}f(r)^{2}h^{\prime}(r)^{2}f^{(4)}(r)-8q_{2}f(r)^{3}h^{\prime}(r)^{2}H^{(4)}(r)$
$+\frac{e^{-2H(r)}f(r)^{2}}{\mathit{k}r^{2}}=0,$ (A-1)
The $rr$-component of the field equations of motion is as follows,
$-112e^{2H(r)}q_{2}\mathit{k}f(r)^{2}h^{\prime}(r)^{2}H^{\prime}(r)^{4}r^{4}+240e^{2H(r)}q_{2}\mathit{k}f(r)f^{\prime}(r)h^{\prime}(r)^{2}H^{\prime}(r)^{3}r^{4}+e^{2H(r)}q_{1}\mathit{k}h^{\prime}(r)^{2}r^{4}$
$-140e^{2H(r)}q_{2}\mathit{k}f^{\prime}(r)^{2}h^{\prime}(r)^{2}H^{\prime}(r)^{2}r^{4}-12e^{2H(r)}q_{2}\mathit{k}h^{\prime}(r)^{2}f^{\prime\prime}(r)^{2}r^{4}-48e^{2H(r)}q_{2}\mathit{k}f(r)^{2}h^{\prime}(r)^{2}H^{\prime\prime}(r)^{2}r^{4}$
$+4\Lambda r^{4}-32e^{2H(r)}q_{2}\mathit{k}f(r)h^{\prime}(r)^{2}H^{\prime}(r)^{2}f^{\prime\prime}(r)r^{4}+64e^{2H(r)}q_{2}\mathit{k}f^{\prime}(r)h^{\prime}(r)^{2}H^{\prime}(r)f^{\prime\prime}(r)r^{4}$
$-64e^{2H(r)}q_{2}\mathit{k}f(r)^{2}h^{\prime}(r)H^{\prime}(r)^{3}h^{\prime\prime}(r)r^{4}+128e^{2H(r)}q_{2}\mathit{k}f(r)f^{\prime}(r)h^{\prime}(r)H^{\prime}(r)^{2}h^{\prime\prime}(r)r^{4}$
$-48e^{2H(r)}q_{2}\mathit{k}f^{\prime}(r)^{2}h^{\prime}(r)H^{\prime}(r)h^{\prime\prime}(r)r^{4}+16e^{2H(r)}q_{2}\mathit{k}f^{\prime}(r)h^{\prime}(r)f^{\prime\prime}(r)h^{\prime\prime}(r)r^{4}$
$-32e^{2H(r)}q_{2}\mathit{k}f(r)h^{\prime}(r)H^{\prime}(r)f^{\prime\prime}(r)h^{\prime\prime}(r)r^{4}-40e^{2H(r)}q_{2}\mathit{k}f^{\prime}(r)^{2}h^{\prime}(r)^{2}H^{\prime\prime}(r)r^{4}$
$+96e^{2H(r)}q_{2}\mathit{k}f(r)^{2}h^{\prime}(r)^{2}H^{\prime}(r)^{2}H^{\prime\prime}(r)r^{4}-64e^{2H(r)}q_{2}\mathit{k}f(r)f^{\prime}(r)h^{\prime}(r)^{2}H^{\prime}(r)H^{\prime\prime}(r)r^{4}$
$+48e^{2H(r)}q_{2}\mathit{k}f(r)h^{\prime}(r)^{2}f^{\prime\prime}(r)H^{\prime\prime}(r)r^{4}-32e^{2H(r)}q_{2}\mathit{k}f(r)f^{\prime}(r)h^{\prime}(r)h^{\prime\prime}(r)H^{\prime\prime}(r)r^{4}$
$+64e^{2H(r)}q_{2}\mathit{k}f(r)^{2}h^{\prime}(r)H^{\prime}(r)h^{\prime\prime}(r)H^{\prime\prime}(r)r^{4}+8e^{2H(r)}q_{2}\mathit{k}f^{\prime}(r)h^{\prime}(r)^{2}f^{(3)}(r)r^{4}$
$-16e^{2H(r)}q_{2}\mathit{k}f(r)h^{\prime}(r)^{2}H^{\prime}(r)f^{(3)}(r)r^{4}-16e^{2H(r)}q_{2}\mathit{k}f(r)f^{\prime}(r)h^{\prime}(r)^{2}H^{(3)}(r)r^{4}$
$+32e^{2H(r)}q_{2}\mathit{k}f(r)^{2}h^{\prime}(r)^{2}H^{\prime}(r)H^{(3)}(r)r^{4}+384e^{2H(r)}q_{2}\mathit{k}f(r)^{2}h^{\prime}(r)^{2}H^{\prime}(r)^{3}r^{3}$
$-608e^{2H(r)}q_{2}\mathit{k}f(r)f^{\prime}(r)h^{\prime}(r)^{2}H^{\prime}(r)^{2}r^{3}+4f^{\prime}(r)r^{3}$
$+272e^{2H(r)}q_{2}\mathit{k}f^{\prime}(r)^{2}h^{\prime}(r)^{2}H^{\prime}(r)r^{3}-8f(r)H^{\prime}(r)r^{3}-48e^{2H(r)}q_{2}\mathit{k}f^{\prime}(r)h^{\prime}(r)^{2}f^{\prime\prime}(r)r^{3}$
$-32e^{2H(r)}q_{2}\mathit{k}f(r)h^{\prime}(r)^{2}H^{\prime}(r)f^{\prime\prime}(r)r^{3}+256e^{2H(r)}q_{2}\mathit{k}f(r)^{2}h^{\prime}(r)H^{\prime}(r)^{2}h^{\prime\prime}(r)r^{3}$
$+64e^{2H(r)}q_{2}\mathit{k}f^{\prime}(r)^{2}h^{\prime}(r)h^{\prime\prime}(r)r^{3}-384e^{2H(r)}q_{2}\mathit{k}f(r)f^{\prime}(r)h^{\prime}(r)H^{\prime}(r)h^{\prime\prime}(r)r^{3}$
$+64e^{2H(r)}q_{2}\mathit{k}f(r)h^{\prime}(r)f^{\prime\prime}(r)h^{\prime\prime}(r)r^{3}-32e^{2H(r)}q_{2}\mathit{k}f(r)f^{\prime}(r)h^{\prime}(r)^{2}H^{\prime\prime}(r)r^{3}$
$-64e^{2H(r)}q_{2}\mathit{k}f(r)^{2}h^{\prime}(r)^{2}H^{\prime}(r)H^{\prime\prime}(r)r^{3}-128e^{2H(r)}q_{2}\mathit{k}f(r)^{2}h^{\prime}(r)h^{\prime\prime}(r)H^{\prime\prime}(r)r^{3}$
$+32e^{2H(r)}q_{2}\mathit{k}f(r)h^{\prime}(r)^{2}f^{(3)}(r)r^{3}-64e^{2H(r)}q_{2}\mathit{k}f(r)^{2}h^{\prime}(r)^{2}H^{(3)}(r)r^{3}$
$-144e^{2H(r)}q_{2}\mathit{k}f^{\prime}(r)^{2}h^{\prime}(r)^{2}r^{2}-512e^{2H(r)}q_{2}\mathit{k}f(r)^{2}h^{\prime}(r)^{2}H^{\prime}(r)^{2}r^{2}+4f(r)r^{2}$
$+512e^{2H(r)}q_{2}\mathit{k}f(r)f^{\prime}(r)h^{\prime}(r)^{2}H^{\prime}(r)r^{2}+96e^{2H(r)}q_{2}\mathit{k}f(r)h^{\prime}(r)^{2}f^{\prime\prime}(r)r^{2}$
$+288e^{2H(r)}q_{2}\mathit{k}f(r)f^{\prime}(r)h^{\prime}(r)h^{\prime\prime}(r)r^{2}-320e^{2H(r)}q_{2}\mathit{k}f(r)^{2}h^{\prime}(r)H^{\prime}(r)h^{\prime\prime}(r)r^{2}$
$-64e^{2H(r)}q_{2}\mathit{k}f(r)^{2}h^{\prime}(r)^{2}H^{\prime\prime}(r)r^{2}-192e^{2H(r)}q_{2}\mathit{k}f(r)f^{\prime}(r)h^{\prime}(r)^{2}r$
$+384e^{2H(r)}q_{2}\mathit{k}f(r)^{2}h^{\prime}(r)^{2}H^{\prime}(r)r+128e^{2H(r)}q_{2}\mathit{k}f(r)^{2}h^{\prime}(r)h^{\prime\prime}(r)r$
$-144e^{2H(r)}q_{2}\mathit{k}f(r)^{2}h^{\prime}(r)^{2}=0.$ (A-2)
Non-zero component of Eq.(10) is as follows,
$-32q_{2}f(r)^{3}h^{\prime}(r)H^{\prime}(r)^{5}r^{5}+32q_{2}f(r)^{2}f^{\prime}(r)h^{\prime}(r)H^{\prime}(r)^{4}r^{5}+24q_{2}f(r)f^{\prime}(r)^{2}h^{\prime}(r)H^{\prime}(r)^{3}r^{5}$
$+40q_{2}f(r)h^{\prime}(r)H^{\prime}(r)f^{\prime\prime}(r)^{2}r^{5}-160q_{2}f(r)^{2}f^{\prime}(r)h^{\prime}(r)H^{\prime\prime}(r)^{2}r^{5}$
$+96q_{2}f(r)^{3}h^{\prime}(r)H^{\prime}(r)H^{\prime\prime}(r)^{2}r^{5}-32q_{2}f(r)^{3}h^{\prime\prime}(r)H^{\prime\prime}(r)^{2}r^{5}+2q_{1}f(r)h^{\prime}(r)H^{\prime}(r)r^{5}$
$+64q_{2}f(r)^{2}h^{\prime}(r)H^{\prime}(r)^{3}f^{\prime\prime}(r)r^{5}-128q_{2}f(r)f^{\prime}(r)h^{\prime}(r)H^{\prime}(r)^{2}f^{\prime\prime}(r)r^{5}$
$-32q_{2}f(r)^{3}H^{\prime}(r)^{4}h^{\prime\prime}(r)r^{5}+96q_{2}f(r)^{2}f^{\prime}(r)H^{\prime}(r)^{3}h^{\prime\prime}(r)r^{5}-72q_{2}f(r)f^{\prime}(r)^{2}H^{\prime}(r)^{2}h^{\prime\prime}(r)r^{5}$
$-8q_{2}f(r)f^{\prime\prime}(r)^{2}h^{\prime\prime}(r)r^{5}+2q_{1}f(r)h^{\prime\prime}(r)r^{5}-32q_{2}f(r)^{2}H^{\prime}(r)^{2}f^{\prime\prime}(r)h^{\prime\prime}(r)r^{5}$
$+48q_{2}f(r)f^{\prime}(r)H^{\prime}(r)f^{\prime\prime}(r)h^{\prime\prime}(r)r^{5}-64q_{2}f(r)^{3}h^{\prime}(r)H^{\prime}(r)^{3}H^{\prime\prime}(r)r^{5}$
$+320q_{2}f(r)^{2}f^{\prime}(r)h^{\prime}(r)H^{\prime}(r)^{2}H^{\prime\prime}(r)r^{5}-240q_{2}f(r)f^{\prime}(r)^{2}h^{\prime}(r)H^{\prime}(r)H^{\prime\prime}(r)r^{5}$
$+80q_{2}f(r)f^{\prime}(r)h^{\prime}(r)f^{\prime\prime}(r)H^{\prime\prime}(r)r^{5}-128q_{2}f(r)^{2}h^{\prime}(r)H^{\prime}(r)f^{\prime\prime}(r)H^{\prime\prime}(r)r^{5}$
$+64q_{2}f(r)^{3}H^{\prime}(r)^{2}h^{\prime\prime}(r)H^{\prime\prime}(r)r^{5}-96q_{2}f(r)^{2}f^{\prime}(r)H^{\prime}(r)h^{\prime\prime}(r)H^{\prime\prime}(r)r^{5}$
$+32q_{2}f(r)^{2}f^{\prime\prime}(r)h^{\prime\prime}(r)H^{\prime\prime}(r)r^{5}-32q_{2}f(r)^{2}h^{\prime}(r)H^{\prime}(r)^{2}f^{(3)}(r)r^{5}$
$+48q_{2}f(r)f^{\prime}(r)h^{\prime}(r)H^{\prime}(r)f^{(3)}(r)r^{5}-16q_{2}f(r)h^{\prime}(r)f^{\prime\prime}(r)f^{(3)}(r)r^{5}$
$+32q_{2}f(r)^{2}h^{\prime}(r)H^{\prime\prime}(r)f^{(3)}(r)r^{5}+64q_{2}f(r)^{3}h^{\prime}(r)H^{\prime}(r)^{2}H^{(3)}(r)r^{5}$
$-96q_{2}f(r)^{2}f^{\prime}(r)h^{\prime}(r)H^{\prime}(r)H^{(3)}(r)r^{5}+32q_{2}f(r)^{2}h^{\prime}(r)f^{\prime\prime}(r)H^{(3)}(r)r^{5}$
$-64q_{2}f(r)^{3}h^{\prime}(r)H^{\prime\prime}(r)H^{(3)}(r)r^{5}+64q_{2}f(r)^{3}h^{\prime}(r)H^{\prime}(r)^{4}r^{4}$
$+128q_{2}f(r)^{2}f^{\prime}(r)h^{\prime}(r)H^{\prime}(r)^{3}r^{4}-272q_{2}f(r)f^{\prime}(r)^{2}h^{\prime}(r)H^{\prime}(r)^{2}r^{4}$
$-80q_{2}f(r)h^{\prime}(r)f^{\prime\prime}(r)^{2}r^{4}-192q_{2}f(r)^{3}h^{\prime}(r)H^{\prime\prime}(r)^{2}r^{4}+4q_{1}f(r)h^{\prime}(r)r^{4}$
$-320q_{2}f(r)^{2}h^{\prime}(r)H^{\prime}(r)^{2}f^{\prime\prime}(r)r^{4}+480q_{2}f(r)f^{\prime}(r)h^{\prime}(r)H^{\prime}(r)f^{\prime\prime}(r)r^{4}$
$+128q_{2}f(r)^{3}H^{\prime}(r)^{3}h^{\prime\prime}(r)r^{4}-320q_{2}f(r)^{2}f^{\prime}(r)H^{\prime}(r)^{2}h^{\prime\prime}(r)r^{4}$
$+192q_{2}f(r)f^{\prime}(r)^{2}H^{\prime}(r)h^{\prime\prime}(r)r^{4}-64q_{2}f(r)f^{\prime}(r)f^{\prime\prime}(r)h^{\prime\prime}(r)r^{4}$
$+64q_{2}f(r)^{2}H^{\prime}(r)f^{\prime\prime}(r)h^{\prime\prime}(r)r^{4}+384q_{2}f(r)^{3}h^{\prime}(r)H^{\prime}(r)^{2}H^{\prime\prime}(r)r^{4}$
$+320q_{2}f(r)f^{\prime}(r)^{2}h^{\prime}(r)H^{\prime\prime}(r)r^{4}-960q_{2}f(r)^{2}f^{\prime}(r)h^{\prime}(r)H^{\prime}(r)H^{\prime\prime}(r)r^{4}$
$+256q_{2}f(r)^{2}h^{\prime}(r)f^{\prime\prime}(r)H^{\prime\prime}(r)r^{4}+128q_{2}f(r)^{2}f^{\prime}(r)h^{\prime\prime}(r)H^{\prime\prime}(r)r^{4}$
$-128q_{2}f(r)^{3}H^{\prime}(r)h^{\prime\prime}(r)H^{\prime\prime}(r)r^{4}-64q_{2}f(r)f^{\prime}(r)h^{\prime}(r)f^{(3)}(r)r^{4}$
$+64q_{2}f(r)^{2}h^{\prime}(r)H^{\prime}(r)f^{(3)}(r)r^{4}+128q_{2}f(r)^{2}f^{\prime}(r)h^{\prime}(r)H^{(3)}(r)r^{4}$
$-128q_{2}f(r)^{3}h^{\prime}(r)H^{\prime}(r)H^{(3)}(r)r^{4}-64q_{2}f(r)^{3}h^{\prime}(r)H^{\prime}(r)^{3}r^{3}$
$-352q_{2}f(r)^{2}f^{\prime}(r)h^{\prime}(r)H^{\prime}(r)^{2}r^{3}+416q_{2}f(r)f^{\prime}(r)^{2}h^{\prime}(r)H^{\prime}(r)r^{3}$
$-352q_{2}f(r)f^{\prime}(r)h^{\prime}(r)f^{\prime\prime}(r)r^{3}+384q_{2}f(r)^{2}h^{\prime}(r)H^{\prime}(r)f^{\prime\prime}(r)r^{3}$
$-128q_{2}f(r)f^{\prime}(r)^{2}h^{\prime\prime}(r)r^{3}-192q_{2}f(r)^{3}H^{\prime}(r)^{2}h^{\prime\prime}(r)r^{3}+352q_{2}f(r)^{2}f^{\prime}(r)H^{\prime}(r)h^{\prime\prime}(r)r^{3}$
$-32q_{2}f(r)^{2}f^{\prime\prime}(r)h^{\prime\prime}(r)r^{3}+608q_{2}f(r)^{2}f^{\prime}(r)h^{\prime}(r)H^{\prime\prime}(r)r^{3}-448q_{2}f(r)^{3}h^{\prime}(r)H^{\prime}(r)H^{\prime\prime}(r)r^{3}$
$+64q_{2}f(r)^{3}h^{\prime\prime}(r)H^{\prime\prime}(r)r^{3}-32q_{2}f(r)^{2}h^{\prime}(r)f^{(3)}(r)r^{3}+64q_{2}f(r)^{3}h^{\prime}(r)H^{(3)}(r)r^{3}$
$+128q_{2}f(r)^{3}h^{\prime}(r)H^{\prime}(r)^{2}r^{2}-128q_{2}f(r)f^{\prime}(r)^{2}h^{\prime}(r)r^{2}+128q_{2}f(r)^{2}f^{\prime}(r)h^{\prime}(r)H^{\prime}(r)r^{2}$
$-128q_{2}f(r)^{2}h^{\prime}(r)f^{\prime\prime}(r)r^{2}-128q_{2}f(r)^{2}f^{\prime}(r)h^{\prime\prime}(r)r^{2}+128q_{2}f(r)^{3}H^{\prime}(r)h^{\prime\prime}(r)r^{2}$
$+128q_{2}f(r)^{3}h^{\prime}(r)H^{\prime\prime}(r)r^{2}+64q_{2}f(r)^{2}f^{\prime}(r)h^{\prime}(r)r$
$-160q_{2}f(r)^{3}h^{\prime}(r)H^{\prime}(r)r-32q_{2}f(r)^{3}h^{\prime\prime}(r)r+64q_{2}f(r)^{3}h^{\prime}(r)=0,$ (A-3)
We write the $tt$-component of Einstein’s equation, Eq. (A), up to first order as,
$-12\mathit{k}r^{4}f_{0}^{\prime\prime}(r)^{2}h_{0}^{\prime}(r)^{2}+16\mathit{k}r^{4}f_{0}^{\prime}(r)f_{0}^{\prime\prime}(r)h_{0}^{\prime}(r)h_{0}^{\prime\prime}(r)+64\mathit{k}r^{4}f_{0}(r)f_{0}^{(3)}(r)h_{0}^{\prime}(r)h_{0}^{\prime\prime}(r)$
$+32\mathit{k}r^{4}f_{0}(r)h_{0}^{(3)}(r)f_{0}^{\prime\prime}(r)h_{0}^{\prime}(r)+128\mathit{k}r^{3}f_{0}(r)f_{0}^{\prime}(r)h_{0}^{\prime\prime}(r)^{2}-48\mathit{k}r^{3}f_{0}^{\prime}(r)f_{0}^{\prime\prime}(r)h_{0}^{\prime}(r)^{2}$
$+64\mathit{k}r^{3}f_{0}^{\prime}(r)^{2}h_{0}^{\prime}(r)h_{0}^{\prime\prime}(r)+320\mathit{k}r^{3}f_{0}(r)f_{0}^{\prime\prime}(r)h_{0}^{\prime}(r)h_{0}^{\prime\prime}(r)+160\mathit{k}r^{2}f_{0}(r)f_{0}^{\prime}(r)h_{0}^{\prime}(r)h_{0}^{\prime\prime}(r)$
$+8\mathit{k}r^{4}f_{0}^{(3)}(r)f_{0}^{\prime}(r)h_{0}^{\prime}(r)^{2}+16\mathit{k}r^{4}f_{0}(r)f_{0}^{(4)}(r)h_{0}^{\prime}(r)^{2}+96\mathit{k}r^{3}f_{0}(r)f_{0}^{(3)}(r)h_{0}^{\prime}(r)^{2}$
$+128\mathit{k}r^{3}f_{0}(r)h_{0}^{(3)}(r)f_{0}^{\prime}(r)h_{0}^{\prime}(r)-144\mathit{k}r^{2}f_{0}^{\prime}(r)^{2}h_{0}^{\prime}(r)^{2}-192\mathit{k}rf_{0}(r)f_{0}^{\prime}(r)h_{0}^{\prime}(r)^{2}$
$+32\mathit{k}r^{4}f_{0}(r)f_{0}^{\prime\prime}(r)h_{0}^{\prime\prime}(r)^{2}-128\mathit{k}rf_{0}(r)^{2}h_{0}^{\prime}(r)h_{0}^{\prime\prime}(r)+64\mathit{k}r^{2}f_{0}(r)^{2}h_{0}^{(3)}(r)h_{0}^{\prime}(r)$
$+48\mathit{k}f_{0}(r)^{2}h_{0}^{\prime}(r)^{2}+64\mathit{k}r^{2}f_{0}(r)^{2}h_{0}^{\prime\prime}(r)^{2}+2r^{3}f_{1}^{\prime}(r)+2r^{2}f_{1}(r)+2\mathit{k}q_{1}r^{4}h_{0}^{\prime}(r)h_{1}^{\prime}(r)$
$+2\mathit{k}q_{1}r^{4}H_{1}(r)h_{0}^{\prime}(r)^{2}$ (A-4)
We write the $rr$-component of Einstein’s equation, Eq. (A), up to first order as,
$-12\mathit{k}r^{4}f_{0}^{\prime\prime}(r)^{2}h_{0}^{\prime}(r)^{2}+16\mathit{k}r^{4}f_{0}^{\prime}(r)f_{0}^{\prime\prime}(r)h_{0}^{\prime}(r)h_{0}^{\prime\prime}(r)-48\mathit{k}r^{3}f_{0}^{\prime}(r)f_{0}^{\prime\prime}(r)h_{0}^{\prime}(r)^{2}$
$+64\mathit{k}r^{3}f_{0}^{\prime}(r)^{2}h_{0}^{\prime}(r)h_{0}^{\prime\prime}(r)+64\mathit{k}r^{3}f_{0}(r)f_{0}^{\prime\prime}(r)h_{0}^{\prime}(r)h_{0}^{\prime\prime}(r)+96\mathit{k}r^{2}f_{0}(r)f_{0}^{\prime\prime}(r)h_{0}^{\prime}(r)^{2}$
$+288\mathit{k}r^{2}f_{0}(r)f_{0}^{\prime}(r)h_{0}^{\prime}(r)h_{0}^{\prime\prime}(r)+8\mathit{k}r^{4}f_{0}^{(3)}(r)f_{0}^{\prime}(r)h_{0}^{\prime}(r)^{2}+32\mathit{k}r^{3}f_{0}(r)f_{0}^{(3)}(r)h_{0}^{\prime}(r)^{2}$
$-144\mathit{k}r^{2}f_{0}^{\prime}(r)^{2}h_{0}^{\prime}(r)^{2}-192\mathit{k}rf_{0}(r)f_{0}^{\prime}(r)h_{0}^{\prime}(r)^{2}+128\mathit{k}rf_{0}(r)^{2}h_{0}^{\prime}(r)h_{0}^{\prime\prime}(r)$
$-144\mathit{k}f_{0}(r)^{2}h_{0}^{\prime}(r)^{2}-4r^{3}f_{0}(r)H_{1}^{\prime}(r)+2r^{3}f_{1}^{\prime}(r)+2r^{2}f_{1}(r)$
$+2\mathit{k}q_{1}r^{4}h_{0}^{\prime}(r)h_{1}^{\prime}(r)+2\mathit{k}q_{1}r^{4}H_{1}(r)h_{0}^{\prime}(r)^{2}$ (A-5)
We impose the condition that Eq. (A) and Eq. (A) must be equivalent up to the first order of $q_{2}$.
$-64\mathit{k}r^{4}f_{0}(r)f_{0}^{(3)}(r)h_{0}^{\prime}(r)h_{0}^{\prime\prime}(r)-32\mathit{k}r^{4}f_{0}(r)h_{0}^{(3)}(r)f_{0}^{\prime\prime}(r)h_{0}^{\prime}(r)-128\mathit{k}r^{3}f_{0}(r)f_{0}^{\prime}(r)h_{0}^{\prime\prime}(r)^{2}$
$-256\mathit{k}r^{3}f_{0}(r)f_{0}^{\prime\prime}(r)h_{0}^{\prime}(r)h_{0}^{\prime\prime}(r)+96\mathit{k}r^{2}f_{0}(r)f_{0}^{\prime\prime}(r)h_{0}^{\prime}(r)^{2}+128\mathit{k}r^{2}f_{0}(r)f_{0}^{\prime}(r)h_{0}^{\prime}(r)h_{0}^{\prime\prime}(r)$
$-16\mathit{k}r^{4}f_{0}(r)f_{0}^{(4)}(r)h_{0}^{\prime}(r)^{2}-64\mathit{k}r^{3}f_{0}(r)f_{0}^{(3)}(r)h_{0}^{\prime}(r)^{2}-128\mathit{k}r^{3}f_{0}(r)h_{0}^{(3)}(r)f_{0}^{\prime}(r)h_{0}^{\prime}(r)$
$-32\mathit{k}r^{4}f_{0}(r)f_{0}^{\prime\prime}(r)h_{0}^{\prime\prime}(r)^{2}+256\mathit{k}rf_{0}(r)^{2}h_{0}^{\prime}(r)h_{0}^{\prime\prime}(r)-64\mathit{k}r^{2}f_{0}(r)^{2}h_{0}^{(3)}(r)h_{0}^{\prime}(r)$
$-192\mathit{k}f_{0}(r)^{2}h_{0}^{\prime}(r)^{2}-64\mathit{k}r^{2}f_{0}(r)^{2}h_{0}^{\prime\prime}(r)^{2}-4r^{3}f_{0}(r)H_{1}^{\prime}(r)=0,$ (A-6)
Appendix B Scrambling time
To compute the scrambling time we need to find the butterfly velocity. In the presence of a particle with momentum $P$ the energy-momentum tensor due to backreaction would be [52]:
$T_{UU}\sim Pe^{\frac{2\pi t}{\beta}}\delta(U)\delta(x)-2K(x,t)\delta(U)T^{0}_{UV},$ (B-7)
$T^{0}_{UV}$ is the energy-momentum without the shock wave, where the backreacted geometry in the Kruskal coordinates would be:
$ds^{2}=2M(UV)dUdV+N(UV)dx^{i}dx^{j}-2M(UV)K(x,t)\delta(U)dU^{2}.$ (B-8)
By plugging the above metric in the Einstein equation:
$(\nabla_{i}-\lambda^{2})K\sim\frac{N(0)Pe^{\frac{2\pi t}{\beta}}\delta(x)}{M(0)},$ (B-9)
where $\lambda^{2}=\frac{d}{2}\frac{N^{\prime}(0)}{M(0)}$ is the effective mass. The solution of $K$ function would be:
$K(x,t)=\frac{e^{\frac{2\pi(t-t_{*})}{\beta}-\lambda x}}{|x|^{\frac{d-1}{2}}}$ (B-10)
Then, the scrambling time would be:
$v_{B}=\frac{2\pi}{\beta\lambda}$ (B-11)
For our solution (12) the butterfly velocity would be:
$v^{2}_{B}$ $\sim\frac{Te^{-H(R)}}{R}=\frac{f^{\prime}(R)e^{-2H(R)}}{4\pi R}=\frac{e^{\frac{2q_{2}Q^{2}}{R^{4}}}}{4\pi R}f^{\prime}(R)$ (B-12)
$\sim 1-Q^{2}(1+q_{2})$ (B-13)
Finally, the butterfly velocity is given by (setting the horizon as unity):
$v_{B}\sim\sqrt{1-Q^{2}(1+q_{2})}$ (B-14)
Then, the scrambling time would be [53]:
$t_{*}(x)\approx\frac{1}{\lambda_{L}}\Big(\log S+\frac{x\lambda_{L}}{v_{B}}\Big),$ (B-15)
$S\sim x^{2},$ (B-16)
where the Lyapunov exponent is $\lambda_{L}=\frac{\beta}{2\pi}$ and $x$ is the size of the system. The butterfly velocity in our solution is small, then the dominant term in the scrambling time is $\frac{x}{v_{B}}$:
$t_{*}\sim\frac{1}{\sqrt{1-Q^{2}(1+q_{2})}}.$ (B-17)